OpenStudy (jinx.exe):

The area of the base of a prism is 55 mm2. The perimeter of the base is 30 mm. The height of the prism is 7 mm. What is the surface area of the prism? A. 445 mm2 B. 320 mm2 C. 265 mm2 D. 92 mm2

OpenStudy (anonymous):

SA = 2(lw + lh + wh)

OpenStudy (anonymous):

area = l*w

OpenStudy (anonymous):

so 7* something gets you 55 but that it also has to be 7 +7 + something + something to get30

OpenStudy (jinx.exe):

wats teh answer doe

OpenStudy (anonymous):

7* what = 55 ?

OpenStudy (anonymous):

nm i did it wrong

OpenStudy (jinx.exe):

omg smh

OpenStudy (anonymous):

here x+y=30:perimeter , xy=55:area

OpenStudy (anonymous):

do some equation like dis y=30−x x(30-x)=55

OpenStudy (anonymous):

ok i tried to help you i cant just give answers

OpenStudy (mathstudent55):

The surface area of a prism can be calculated several ways. One way is to use the area of the bases, the perimeter of the base and the heioght, which is exactly what you have. \(SA = 2 \times ~area~of~base + perimeter~of~base \times height\) \(SA = 2B + Ph\) where SA = surface area B = area of one base P = perimeter of base h = height of prism